Jake and Bob scored a total of 16 touchdowns. Bob scored one more than twice the number of Jake's touchdowns. How many did each score?
Answer by jim_thompson5910(35256)

Let x = # of touchdowns Jake scored and y = # of touchdowns Bob scored

x+y = 16
y = 2x+1


x+y = 16

x+2x+1 = 16

3x+1 = 16

3x = 16-1

3x = 15

x = 15/3

x = 5

y = 2x+1

y = 2(5)+1

y = 10+1

y = 11

Answer:

Jake scored 5 touchdowns.

Bob scored 11 touchdowns.